# DecAlign: Hierarchical Cross-Modal Alignment for Decoupled Multimodal Representation Learning

### v1

@@ 3.3 Homogeneity Alignment
While different modalities exhibit unique characteristics in their representations, they also share common elements that convey the same semantic information. To effectively uncovering and aligning these shared features, it is crucial to address the inherent challenges posed by modality-unique variations and residual inconsistencies in their distributions.
3.3.1 Latent Space Semantic Alignment
To address the global offset and semantic inconsistencies in modality-common features and mitigate information distortion during feature fusion, we model modality feature distributions using Gaussian distributions. By mapping representations into a latent space, we quantify differences in position, shape, and symmetry through mean, covariance, and skewness. Specifically, for modality-common features, their distributions are approximated as $\mathcal{Z}^{m_{i}}_{com}\sim\mathcal{N}(\mu_{com}^{m_{i}},\Sigma_{com}^{m_{i}},\Gamma_{com}^{m_{i}})$, where $\mu_{com}^{m_{i}}$, $\Sigma_{com}^{m_{i}}$ and $\Gamma_{com}^{m_{i}}$ represent the mean, covariance and skewness of the common features for modality $m_{i}$, respectively. Their detailed formulas are discussed in the Appendix. To ensure semantic consistency across modalities, we define the latent space semantic alignment loss as:
$\begin{split}\mathcal{L}_{sem}&=\frac{1}{M(M-1)}\sum_{1\leq i<j\leq M}\Big{(}||\mu_{com}^{m_{i}}-\mu_{com}^{m_{j}}||^{2}\\ &+||\Sigma_{com}^{m_{i}}-\Sigma_{com}^{m_{j}}||^{2}_{F}+||\Gamma_{com}^{m_{i}}-\Gamma_{com}^{m_{j}}||^{2}\Big{)}\end{split}$ (10)
3.3.2 Cross-Modal Distribution Alignment
To flexibly model the latent distribution space of modality-homogeneous features extracted by the shared encoder without relying on prior knowledge, we employ Maximum Mean Discrepancy (MMD), which maps data to a high-dimensional space using kernel functions and computes the distance between their means, enabling non-parametric and distribution-agnostic modeling. The discrepancy of cross-modal distribution is then quantified as:
$\begin{split}\mathcal{L}_{\text{MMD}}&=\frac{2}{M(M-1)}\sum_{1\leq i<j\leq M}\Big{[}\mathbb{E}_{x,x^{\prime}\sim\mathcal{Z}_{com}^{m_{i}}}[k(x,x^{\prime})]\\ &+\mathbb{E}_{y,y^{\prime}\sim\mathcal{Z}_{com}^{m_{j}}}[k(y,y^{\prime})]-2~{}\mathbb{E}_{x\sim\mathcal{Z}_{com}^{m_{i}},y\sim\mathcal{Z}_{com}^{m_{j}}}[k(x,y)]\Big{]}\end{split}$ (11)
@@ 3.4 Multimodal Fusion and Prediction
Recognizing the unique characteristics of multimodal heterogeneous representations, such as syntactic structures in language, spatial relationships in vision, and temporal patterns in audio, we incorporate multimodal transformer [35] to model global temporal information, providing finer-grained and more consistent feature representations across modalities. These features are then fused with modality-common features via a concatenation operation and passed through a fully connected layer to produce the final prediction results. The overall optimization objective of our framework is defined as:
@@
where $\mathcal{L}_{task}$ represents the task-specific loss, such as cross-entropy for classification tasks or mean squared error for regression. $\alpha$ and $\beta$ are trade-off hyper-parameters for the losses of heterogeneous and homogeneous alignment, with their sensitivity analyzed in Section 4.3.
Models CMU-MOSI CMU-MOSEI
MAE ($\downarrow$) Corr ($\uparrow$) Acc-2 ($\uparrow$) Acc-7 ($\uparrow$) F1 Score ($\uparrow$) MAE ($\downarrow$) Corr ($\uparrow$) Acc-2 ($\uparrow$) Acc-7 ($\uparrow$) F1 Score ($\uparrow$)
MFM [34] 0.951 0.662 78.18 36.21 78.10 0.681 0.555 78.93 45.93 76.45
MulT [35] 0.846 0.725 81.70 40.05 81.66 0.673 0.677 80.85 48.37 80.86
PMR [6] 0.895 0.689 79.88 40.60 79.83 0.645 0.689 81.57 48.88 81.56
CubeMLP [31] 0.838 0.695 81.85 41.03 81.74 0.601 0.701 81.36 49.07 81.75
MUTA-Net [32] 0.767 0.736 82.12 40.88 82.07 0.617 0.717 81.76 49.88 82.01
MISA [11] 0.788 0.744 82.07 41.27 82.43 0.594 0.724 82.03 51.43 82.13
CENet [36] 0.745 0.749 82.40 41.32 82.56 0.588 0.738 82.13 52.31 82.35
Self-MM [50] 0.765 0.764 82.88 42.03 83.04 0.576 0.732 82.43 52.68 82.47
FDMER [47] 0.760 0.777 83.01 42.88 83.22 0.571 0.743 83.88 53.21 83.35
AOBERT [14] 0.780 0.773 83.03 43.21 83.02 0.588 0.738 83.90 52.47 83.14
DMD [17] 0.744 0.788 83.24 43.88 83.55 0.561 0.758 84.17 54.18 83.88
MCIS [48] 0.756 0.783 84.02 43.58 83.85 0.557 0.747 84.97 53.85 84.34
CGGM [9] 0.747 0.798 84.43 43.21 84.13 0.551 0.761 83.90 53.47 84.14
DecAlign (Ours) 0.735 0.811 85.75 45.07 85.82 0.543 0.768 86.48 55.02 86.07
Table 1: Performance Comparison on CMU-MOSI and CMU-MOSEI datasets. $\uparrow$ and $\downarrow$ indicate that higher or lower value is better. Best results are highlighted in bold, and suboptimal results are underlined.
@@ 4 Experiments
Datasets. To ensure direct and fair comparison with prior works, we evaluate DecAlign on four widely used benchmarking datasets, all of which contain textual, visual, and acoustic modalities. CMU-MOSI [51] (2199 clips) and CMU-MOSEI [52] (22856 clips) provide word-aligned multimodal signals. Each sample is annotated with a sentiment score ranging from -3 (highly negative) to +3 (highly positive). CH-SIMS [49] includes 38,280 Chinese utterances with sentiment score ranging from -1 (negative) to +1 (positive). IEMOCAP [3] features 10,039 dynamic utterances annotated for six emotions: angry, happy, sad, neutral, excited, and frustrated.
Evaluation Metrics. For CMU-MOSI and CMU-MOSEI, following prior works [21, 17, 57], we evaluate performance using binary accuracy (Acc-2), 7-class accuracy (Acc-7), and Binary F1 Score. Acc-2 reflects whether the a sample is predicted as negative, while sentiment intensity prediction is further assessed via Mean Absolute Error (MAE) and Pearson Correlation (Corr) to capture deviation and linearity. For CH-SIMS, we adopt MAE and F1 Score. IEMOCAP follows [18, 7, 55] with weighted accuracy (WAcc) and weighted average F1 Score (WAF1), accounting for class distribution imbalance.
Implementation Details. Consistent with previous studies [17, 39], we apply MMSA-FET Toolkit [49, 50] to extract features except for IEMOCAP dataset. For IEMOCAP, we follow prior work [18] for pre-processing. We adopt Adam as optimizer and implement all the experiments on an NVIDIA A6000 ADA with 48 GB memory. Pre-processing details and configuration settings are listed in Appendix.
Models CH-SIMS IEMOCAP (six-class)
MAE ($\downarrow$) F1 Score ($\uparrow$) WAcc ($\uparrow$) WAF1 ($\uparrow$)
MFM [34] 0.471 75.28 63.38 63.41
MulT [35] 0.455 76.96 65.53 65.21
PMR [6] 0.445 76.55 67.04 67.01
CubeMLP [31] 0.459 77.85 66.43 66.41
MUTA-Net [32] 0.443 77.21 67.44 68.78
MISA [11] 0.437 78.43 68.48 68.25
CENet [36] 0.454 78.03 69.27 69.58
Self-MM [50] 0.432 77.97 70.35 70.43
FDMER [47] 0.424 78.74 71.33 71.17
AOBERT [14] 0.430 78.55 71.04 70.89
DMD [17] 0.421 79.88 72.03 71.98
MCIS [48] 0.429 79.58 71.77 71.53
CGGM [9] 0.417 80.12 72.25 72.17
DecAlign (Ours) 0.403 81.85 73.35 73.43
Table 2: Performance comparison on CH-SIMS and IEMOCAP datasets. $\uparrow$ indicates that higher values are better. Best results are highlighted in bold, and runner-up results are underlined.
(a) MulT [35]
(b) MISA [11]
(c) DMD [17]
@@
Figure 4: Confusion matrix comparison with state-of-the-art methods, illustrating the distribution of predicted sentiment categories against ground truth labels on the CMU-MOSI dataset.
4.1 Comparison with State-of-the-art Methods
We compare our DecAlign approach with state-of-the-art methods under unified environmental and dataset settings: MFM [34], MulT [35], PMR [6], CubeMLP [31], MUTA-Net [32], MISA [11], CENet [36], Self-MM [50], FDMER [47], AOBERT [14], DMD [17], MCIS [48], and CGGM [9]. Table 1 & 2, along with Figure 2, presents a comprehensive comparison of our proposed DecAlign framework against state-of-the-art methods on four datasets, evaluated across five metrics which jointly assess the model’s performance from both regression and classification perspectives. The results demonstrate that DecAlign consistently outperforms all baselines, achieving the lowest MAE and highest Corr, indicating superior performance in sentiment intensity estimation, as well as the highest classification accuracy across Acc-2, Acc-7, and F1 Score, confirming its ability to capture both fine-grained and high-level semantic patterns.
@@ 4.1 Comparison with State-of-the-art Methods
Transformer-based methods. Compared to Transformer-based methods such as MulT [35], Self-MM [50], PMR [6], and MUTA-Net [32], which rely on cross-attention mechanism for global feature fusion, DecAlign overcomes modality-unique interference and local semantic inconsistencies. Transformer-based models assume a shared latent space, often causing dominant modalities to overshadow weaker ones, leading to information loss. In contrast, DecAlign explicitly disentangles modality-heterogeneous and modality-homogeneous features, leveraging prototype-based optimal transport for fine-grained alignment and latent space semantic alignment with MMD regularization for global consistency. This mitigates modality interference, reducing MAE and improving Corr, while enhancing classification performance.
@@
Feature Decoupling-based methods. While multimodal feature decoupling methods such as MISA [11], FDMER [47], and DMD [17] alleviate modality interference, they primarily focus on global alignment, often overlooking token-level inconsistencies. This limitation hinders fine-grained multimodal integration, particularly in tasks requiring precise semantic fusion. DecAlign overcomes this challenge through a dual-stream hierarchical alignment strategy, integrating prototype-based transport for local alignment with semantic consistency constraints for robust global integration. This enables more expressive multimodal representations, leading to superior performance across both regression and classification metrics.
Modules CMU-MOSI CMU-MOSEI
MFD Hete Homo MAE F1 Score MAE F1 Score
0.747 84.46 0.562 84.74
0.754 84.03 0.588 84.37
0.784 81.92 0.632 82.22
0.794 81.56 0.624 81.87
Table 3: Ablation study on different key modules for CMU-MOSI and CMU-MOSEI datasets.
Modules CMU-MOSI CMU-MOSEI
Proto-OT CT Sem MMD MAE F1 Score MAE F1 Score
0.741 84.61 0.564 85.26
0.738 84.73 0.553 85.33
0.743 84.36 0.619 85.21
0.748 84.17 0.624 85.03
Table 4: Ablation study on different alignment strategies for CMU-MOSI and CMU-MOSEI datasets.
@@
Confusion Matrix Analysis. To further demonstrate the superiority of our performance and validate the effectiveness of our proposed approach, we analyze the confusion matrix of DecAlign in comparison with representative works in the field of multimodal sentiment analysis, including MulT [35], MISA [11], and DMD [17]. As shown in Figure 4, DecAlign achieves a more balanced and accurate sentiment classification across different sentiment intensity levels, significantly reducing misidentification errors, particularly in distinguishing subtle sentiment variations.
@@ 4.2 Ablation Studies
(d) IEMOCAP (six-class)
@@
Figure 5: Visualization of ablation studies on accuracy comparison across different emotion categories.
@@
Impact of key components. We evaluate the impact of Multimodal Feature Decoupling (MFD), Heterogeneous (Hete), and Homogeneous (Homo) Alignment on model performance using MAE, Acc-2, and F1 Score (Table 3). The full model achieves the best results, confirming the significance of hierarchical alignment. Removing Homogeneous Alignment slightly increases MAE and lowers Acc-2, indicating the importance of intra-modal consistency. Eliminating Heterogeneous Alignment leads to a greater drop, showing that modality-unique interference affects feature integration. The absence of both alignments causes substantial performance degradation, highlighting the need to disentangle modality-homogeneous and modality-heterogeneous features.
@@
Additionally, Figure 5 visualizes the ablation results across different sentiment categories, illustrating the performance variations when heterogeneous and homogeneous alignment modules are frozen. The degradation across sentiment categories further validates the necessity of a hierarchical alignment strategy to maintain robust performance across diverse emotional expressions. Notably, even when any single alignment module is disabled, the F1 Score remains higher than many state-of-the-art methods, including FDMER, AOBERT, and DMD, demonstrating the effectiveness of our proposed alignment approach from both heterogeneous and homogeneous perspectives. The most severe performance degradation occurs when MFD is removed, demonstrating that preserving modality-unique information before fusion is crucial. This underscores the effectiveness of integrating heterogeneous and homogeneous representations for better sentiment analysis.
@@
Impact of specific alignment strategies. We further evaluate the contribution of Prototype-Based Optimal Transport (Proto-OT), Contrastive Training (CT), Semantic Consistency (Sem), and Maximum Mean Discrepancy (MMD) Regularization to DecAlign’s performance, as shown in Table 4. Removing MMD regularization leads to a slight performance drop, highlighting its role in global latent space alignment and feature coherence. The exclusion of semantic consistency further degrades performance, indicating that enforcing semantic alignment enhances multimodal feature integration. The most substantial drop occurs when contrastive training is removed, showing its critical role in learning discriminative multimodal representations. Similarly, eliminating Proto-OT results in a notable decline in both regression and classification metrics, demonstrating that fine-grained alignment through optimal transport significantly improves multimodal fusion and overall sentiment prediction accuracy.
@@
Analysis of modality Gap. Figure 6 presents a case study on vision and language modalities, demonstrating how DecAlign mitigates the modality gap to enhance alignment. Models without heterogeneous or homogeneous alignment exhibit significantly larger gaps, hindering cross-modal fusion. These results further validate the effectiveness of our hierarchical alignment strategy.
(a) w/o Hete & Homo
(b) w/o Hete
(c) DecAlign (Ours)
Figure 6: Visualization of the modality gap between vision and language on CMU-MOSEI dataset.

### v2

@@ 3.3 Homogeneity Alignment
While different modalities exhibit unique characteristics in their representations, they also share common elements that convey the same semantic information. To effectively uncover and align these shared features, it is crucial to address the inherent challenges posed by modality-unique variations and residual inconsistencies in their distributions.
Latent Space Semantic Alignment. To address the global offset and semantic inconsistencies in modality-common features and mitigate information distortion during feature fusion, we model modality feature distributions using Gaussian distributions. By mapping representations into a latent space, we quantify differences in position, shape, and symmetry through mean, covariance, and skewness, where skewness is further incorporated to capture asymmetry in the distribution of modality-common features, enabling the alignment to account for non-Gaussian semantic variations and improve cross-modal consistency. Specifically, for modality-common features, their distributions are approximated as $\mathcal{Z}^{m_{i}}_{com}\sim\mathcal{N}(\mu_{com}^{m_{i}},\Sigma_{com}^{m_{i}},\Gamma_{com}^{m_{i}})$, where $\mu_{com}^{m_{i}}$, $\Sigma_{com}^{m_{i}}$ and $\Gamma_{com}^{m_{i}}$ represent the mean, covariance and skewness of the common features for modality $m_{i}$, respectively. Their detailed formulas are discussed in the Appendix B.6. To ensure semantic consistency across modalities, we define the latent space semantic alignment loss as:
[…]
Cross-Modal Distribution Alignment. To flexibly model the latent distribution space of modality-homogeneous features extracted by the shared encoder without relying on prior knowledge, we use Probabilistic Distribution Encoder (PDE) to encode feature distributions in latent space. PDE outputs are compared across modalities using the Maximum Mean Discrepancy (MMD) metric, which evaluates the distance between distributions by mapping them into a Reproducing Kernel Hilbert Space (RKHS) and measuring the difference between their mean embeddings. This kernel-based formulation enables non-parametric modeling and captures higher-order statistical properties in a unified space. The discrepancy of cross-modal distribution is then quantified as:
$\begin{split}\mathcal{L}_{\text{MMD}}&=\frac{2}{M(M-1)}\sum_{1\leq i<j\leq M}\Big[\mathbb{E}_{x,x^{\prime}\sim\mathcal{Z}_{com}^{m_{i}}}[k(x,x^{\prime})]\\ &\hskip-18.00005pt+\mathbb{E}_{y,y^{\prime}\sim\mathcal{Z}_{com}^{m_{j}}}[k(y,y^{\prime})]-2~\mathbb{E}_{x\sim\mathcal{Z}_{com}^{m_{i}},y\sim\mathcal{Z}_{com}^{m_{j}}}[k(x,y)]\Big]\end{split}$ (11)
@@ 3.4 Multimodal Fusion and Prediction
Recognizing the unique characteristics of multimodal heterogeneous representations—such as syntactic structures in language, spatial layouts in vision, and temporal patterns in audio—we incorporate modality-specific transformers (Tsai et al., 2019) to enhance global temporal and contextual modeling. While prior alignment places modality-unique features in semantically consistent spaces, these representations still contain rich intra-modal information that benefits from further refinement. Using separate transformers per modality does not undermine alignment, as the representation space has been regularized by alignment losses. Instead, the transformers serve as modality-aware refiners. Their outputs are concatenated with modality-common features, enabling both shared semantics and modality-specific cues to jointly inform the final prediction, which is generated by a fully connected layer. The overall optimization objective of our framework is defined as:
@@
where $\mathcal{L}_{task}$ represents the task-specific loss, such as cross-entropy for classification tasks or mean squared error for regression. $\alpha$ and $\beta$ are trade-off hyperparameters for the losses of heterogeneous and homogeneous alignment, with their sensitivity analyzed in Section 4.3.
@@ 4 Experiments
Dataset and Metric Description. We evaluate DecAlign on four common multimodal datasets: CMU-MOSI (Zadeh et al., 2016), CMU-MOSEI (Zadeh et al., 2018), CH-SIMS (Yu et al., 2020) and IEMOCAP (Busso et al., 2008). For CMU-MOSI and CMU-MOSEI, following prior works (Liang et al., 2021; Li et al., 2023; Zhou et al., 2025), we evaluate performance using binary accuracy (Acc-2), 7-class accuracy (Acc-7), and Binary F1 Score. Acc-2 reflects whether the a sample is predicted as negative, while sentiment intensity prediction is further assessed via Mean Absolute Error (MAE) and Pearson Correlation (Corr) to capture deviation and linearity. For CH-SIMS, we adopt MAE and F1 Score. IEMOCAP follows (Lian et al., 2023; Fu et al., 2024; Zhang et al., 2024) with weighted accuracy (WAcc) and weighted average F1 Score (WAF1), accounting for class distribution imbalance. Detailed dataset and metric descriptions are provided in Appendix B.
Implementation Details. Consistent with previous studies (Li et al., 2023; Wang et al., 2023), we use the MMSA-FET Toolkit (Yu et al., 2021) for feature extraction on all datasets except IEMOCAP, for which we follow the pre-processing procedure described in prior representative work (Lian et al., 2023). We train DecAlign for 50 epochs using Adam optimizer with a batch size of 32 on an NVIDIA A6000. Further details regarding hyperparameter settings are provided in Appendix B.3, and feature extraction is described in Appendix B.4.
4.1 Comparison Analysis (Extended Version in Appendix C)
We compare DecAlign with a range of state-of-the-art methods under a unified experimental environment and consistent dataset splits. These baselines include MFM (Tsai et al., 2018), MulT (Tsai et al., 2019), PMR (Fan et al., 2023), CubeMLP (Sun et al., 2022), MUTA-Net (Tang et al., 2023), MISA (Hazarika et al., 2020), CENet (Wang et al., 2022a), Self-MM (Yu et al., 2021), FDMER (Yang et al., 2022), AOBERT (Kim & Park, 2023), DMD (Li et al., 2023), ReconBoost (Hua et al., 2024), and CGGM (Guo et al., 2025). Table 1567, along with Figure 1, present a comprehensive comparison of our DecAlign framework against 13 state-of-the-art methods on four widely used datasets. To account for statistical significance and reduce the influence of randomness, the reported performance of DecAlign is averaged over five independent runs. The comparison reveals that DecAlign exhibits a stronger ability to capture subtle variations in continuous target values, as well as a more precise distinction among discrete categories. Its consistent performance across diverse datasets indicates an enhanced capacity for modeling both continuous and categorical patterns within multimodal data, reflecting a more comprehensive understanding of complex cross-modal interactions.
(a) MulT
(b) MISA
(c) DMD
@@ 4.1 Comparison Analysis (Extended Version in Appendix C)
Figure 3: Comparison of predicted versus ground truth category distributions for four representative models on the CMU-MOSI dataset.
Models CMU-MOSI CMU-MOSEI IEMOCAP (six-class) CH-SIMS
MAE ($\downarrow$) Acc-2 ($\uparrow$) F1 Score ($\uparrow$) MAE ($\downarrow$) Acc-2 ($\uparrow$) F1 Score ($\uparrow$) WAcc ($\uparrow$) WAF1 ($\uparrow$) MAE ($\downarrow$) F1 Score ($\uparrow$)
MFM (Tsai et al., 2018) 0.951 78.18 78.10 0.681 78.93 76.45 63.38 63.41 0.471 75.28
MulT (Tsai et al., 2019) 0.846 81.70 81.66 0.673 80.85 80.86 65.53 65.21 0.455 76.96
PMR (Fan et al., 2023) 0.895 79.88 79.83 0.645 81.57 81.56 67.04 67.01 0.445 76.55
CubeMLP (Sun et al., 2022) 0.838 81.85 81.74 0.601 81.36 81.75 66.43 66.41 0.459 77.85
MUTA-Net (Tang et al., 2023) 0.767 82.12 82.07 0.617 81.76 82.01 67.44 68.78 0.443 77.21
MISA (Hazarika et al., 2020) 0.788 82.07 82.43 0.594 82.03 82.13 68.48 68.25 0.437 78.43
CENet (Wang et al., 2022a) 0.745 82.40 82.56 0.588 82.13 82.35 69.27 69.58 0.454 78.03
Self-MM (Yu et al., 2021) 0.765 82.88 83.04 0.576 82.43 82.47 70.35 70.43 0.432 77.97
FDMER (Yang et al., 2022) 0.760 83.01 83.22 0.571 83.88 83.35 71.33 71.17 0.424 78.74
AOBERT (Kim & Park, 2023) 0.780 83.03 83.02 0.588 83.90 83.64 71.04 70.89 0.430 78.55
DMD (Li et al., 2023) 0.744 83.24 83.55 0.561 84.17 83.88 72.03 71.98 0.421 79.88
ReconBoost (Hua et al., 2024) 0.793 82.59 82.72 0.599 82.98 83.14 71.44 71.58 0.413 80.41
CGGM (Guo et al., 2025) 0.787 82.73 82.89 0.584 83.72 83.94 72.25 72.17 0.417 80.12
DecAlign (Ours) 0.735 85.75 85.82 0.543 86.48 86.07 73.35 73.43 0.403 81.85
Table 1: Performance comparison across four widely used datasets under a unified experimental setting with consistent data splits to ensure a fair evaluation. Symbols $\uparrow$ and $\downarrow$ indicate that higher or lower values are better, respectively. Best results are highlighted in bold, and second-best results are underlined. All reported results are averaged over five runs on the test set.
@@
Transformer-based methods. Compared to Transformer-based methods such as MulT (Tsai et al., 2019), Self-MM (Yu et al., 2021), PMR (Fan et al., 2023), and MUTA-Net (Tang et al., 2023), which rely on cross-attention mechanism for global feature fusion, DecAlign overcomes modality-unique interference and local semantic inconsistencies. Transformer-based models assume a shared latent space, often causing dominant modalities to overshadow weaker ones, leading to information loss. In contrast, DecAlign explicitly disentangles modality-heterogeneous and modality-homogeneous features, leveraging prototype-based optimal transport for fine-grained alignment and latent space semantic alignment with MMD regularization for global consistency. This mitigates modality interference, reducing MAE and improving Corr, while enhancing classification performance.
@@
Feature Decoupling-based methods. While multimodal feature decoupling methods such as MISA (Hazarika et al., 2020), FDMER (Yang et al., 2022), and DMD (Li et al., 2023) alleviate modality interference, they primarily focus on global alignment, often overlooking token-level inconsistencies. This limitation hinders fine-grained multimodal integration, particularly in tasks requiring precise semantic fusion. DecAlign overcomes this challenge through a dual-stream hierarchical alignment strategy, integrating prototype-based transport for local alignment with semantic consistency constraints for robust global integration. This enables more expressive multimodal representations, leading to superior performance across both regression and classification metrics.
@@
Confusion Matrix Analysis. To further demonstrate the superiority of our performance and validate the effectiveness of our proposed approach, we analyze the confusion matrix of DecAlign in comparison with representative works in the field of multimodal sentiment analysis, including MulT (Tsai et al., 2019), MISA (Hazarika et al., 2020), and DMD (Li et al., 2023). As shown in Figure 3, DecAlign achieves a more balanced and accurate sentiment classification across different sentiment intensity levels, significantly reducing misidentification errors, particularly in distinguishing subtle sentiment variations.
@@
(d) IEMOCAP
(e) w/o Hete & Homo
(f) w/o Hete
(g) w/o Homo
(h) DecAlign (Ours)
@@
Figure 4: Visualization of Ablation Studies. (a)–(d) illustrate the performance comparison across different emotion categories on four benchmarks, (e)–(h) visualize the modality gap between visual and language modalities on the CMU-MOSEI dataset.
Key Modules CMU-MOSI CMU-MOSEI Alignment Strategies CMU-MOSI CMU-MOSEI
MFD Hete Homo MAE F1 MAE F1 Proto-OT CT Sem MMD MAE F1 MAE F1
0.747 84.46 0.562 84.74 0.741 84.61 0.564 85.26
0.754 84.03 0.588 84.37 0.738 84.73 0.553 85.33
0.784 81.92 0.632 82.22 0.743 84.36 0.619 85.21
0.794 81.56 0.624 81.87 0.748 84.17 0.624 85.03
Table 2: Ablation study on key modules (left) and alignment strategies (right) for CMU-MOSI and CMU-MOSEI datasets.
4.2 Ablation Studies (Extended Version in Appendix C.3)
@@
Impact of key components. We evaluate the impact of Multimodal Feature Decoupling (MFD), Heterogeneous (Hete), and Homogeneous (Homo) Alignment on model performance using MAE and Binary F1 Score (Table 2). The full model achieves the best results, confirming the significance of hierarchical alignment. Removing Homogeneous Alignment slightly increases MAE and lowers Acc-2, indicating the importance of intra-modal consistency. Eliminating Heterogeneous Alignment leads to a greater drop, showing that modality-unique interference affects feature integration. The absence of both alignments causes substantial performance degradation, highlighting the need to disentangle modality-homogeneous and modality-heterogeneous features.
@@
Additionally, Figure 4 (a)-(d) visualizes the ablation results across different sentiment categories, illustrating the performance variations when heterogeneous and homogeneous alignment modules are frozen. The degradation across sentiment categories further validates the necessity of a hierarchical alignment strategy to maintain robust performance across diverse emotional expressions. Notably, even when any single alignment module is disabled, the F1 Score remains higher than many state-of-the-art methods, including FDMER, AOBERT, and DMD, demonstrating the effectiveness of our proposed alignment approach from both heterogeneous and homogeneous perspectives. The most severe performance degradation occurs when MFD is removed, demonstrating that preserving modality-unique information before fusion is crucial. This underscores the effectiveness of integrating heterogeneous and homogeneous representations for better sentiment analysis.
@@
Impact of specific alignment strategies. We further evaluate the contribution of Prototype-Based Optimal Transport (Proto-OT), Contrastive Training (CT), Semantic Consistency (Sem), and Maximum Mean Discrepancy (MMD) Regularization to DecAlign’s performance, as shown in Table 2. Removing MMD regularization leads to a slight performance drop, highlighting its role in global latent space alignment and feature coherence. The exclusion of semantic consistency further degrades performance, indicating that enforcing semantic alignment enhances multimodal feature integration. The most substantial drop occurs when contrastive training is removed, showing its critical role in learning discriminative multimodal representations. Similarly, eliminating Proto-OT results in a notable decline in both regression and classification metrics, demonstrating that fine-grained alignment through optimal transport significantly improves multimodal collaborative prediction performance.
@@
Analysis of modality Gap. Figure 4 (e)-(h) presents a case study on vision and language modalities, demonstrating how DecAlign mitigates the modality gap to enhance alignment. Models without heterogeneous or homogeneous alignment exhibit significantly larger gaps, hindering cross-modal fusion. These results further validate the effectiveness of our hierarchical alignment strategy. Extended Analysis will be shown in Appendix C.4.
4.3 Parameter Sensitivity Analysis